OpenStudy (anonymous):

the variables x and y vary directly. write an equation that relates x and y. then find x when y=-4. x=-20/3, y=-15/8

jimthompson5910 (jim_thompson5910):

the variables x and y vary directly means that y = kx

jimthompson5910 (jim_thompson5910):

y = kx -20/3 = k*(-15/8) ... Plug in x=-20/3, y=-15/8 Now solve for k

OpenStudy (anonymous):

32/9?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

y = kx y = (32/9)*x Now plug in y = -4 and solve for x -4 = (32/9)*x keep going to solve for x

OpenStudy (anonymous):

-8/9?

jimthompson5910 (jim_thompson5910):

close, but no

jimthompson5910 (jim_thompson5910):

double check your work

OpenStudy (anonymous):

i still got the same answer..

jimthompson5910 (jim_thompson5910):

-4 = (32/9)*x -4*9 = 32x -36 = 32x did you get this far?

OpenStudy (anonymous):

yep. -9/8?

jimthompson5910 (jim_thompson5910):

good, -9/8 is the answer